OpenStudy (anonymous):

Find the limit as x approaches infinity of sqrt(x^4+4x^3+5) / 10x^2-8

OpenStudy (raden):

hint : the numerator and the denominator divided by x^2

OpenStudy (anonymous):

at @RadEn so I did that and I can see that if I look at the exponents and since its x^6/x^4 I would get my limit to be 1/10 ?

OpenStudy (raden):

the answer is right, it is 1/10. but i dont knw where did you get x^6/x^4 ?

OpenStudy (anonymous):

Well the x^6/x^4 was just part of my answer, sorry I did not type my work. I hadn't thought of dividing by x^2, so thank you so much for your hint :)

OpenStudy (raden):

actually, if the numerator is divided by x^2, get sqrt(x^4+4x^3+5)/x^2 = sqrt(x^4+4x^3+5)/sqrt(x^4) = sqrt((x^4+4x^3+5)/x^4) = sqrt(1+4/x+5/x^4) and if the denominator divided by x^2, get (10x^2-8 )/x^2 = 10 - 8/x^2 so, the limit becomes limit x->inf sqrt(1+4/x+5/x^4)/(10 - 8/x^2) = 1/10

OpenStudy (anonymous):

Oh I just looked at my work and I did some mistakes but now I understand it better, thank you.

OpenStudy (raden):

you're welcome